# Geolocation Representation from Large Language Models are Generic Enhancers for Spatio-Temporal Learning

### v1

@@ Abstract
In the geospatial domain, universal representation models are significantly less prevalent than their extensive use in natural language processing and computer vision. This discrepancy arises primarily from the high costs associated with the input of existing representation models, which often require street views and mobility data. To address this, we develop a novel, training-free method that leverages large language models (LLMs) and auxiliary map data from OpenStreetMap to derive geolocation representations (LLMGeovec). LLMGeovec can represent the geographic semantics of city, country, and global scales, which acts as a generic enhancer for spatio-temporal learning. Specifically, by direct feature concatenation, we introduce a simple yet effective paradigm for enhancing multiple spatio-temporal tasks including geographic prediction (GP), long-term time series forecasting (LTSF), and graph-based spatio-temporal forecasting (GSTF). LLMGeovec can seamlessly integrate into a wide spectrum of spatio-temporal learning models, providing immediate enhancements. Experimental results demonstrate that LLMGeovec achieves global coverage and significantly boosts the performance of leading GP, LTSF, and GSTF models.
@@ Introduction
The reason why previous approaches do not further extend the complex applications is that they do not achieve global coverage and they have a heavy reliance on expensive input data such as street views, travel patterns, and traffic trajectories  (Wang, Li, and Rajagopal 2020; Kim and Yoon 2022; Lee et al. 2021; Zhang et al. 2023a). Although studies have used free and globally available satellite imagery for geolocation representation, their effectiveness has been hampered by the low resolution and absence of important features such as activity patterns  (Manvi et al. 2023; Robinson, Hohman, and Dilkina 2017; Head et al. 2017; Jean et al. 2019; Elmustafa et al. 2022; Xi et al. 2022).
@@
In this paper, we introduce a novel, training-free method that utilizes LLMs and auxiliary map data from OpenStreetMap to derive geolocation representations (LLMGeovec). As illustrated in Fig.  1, our approach first extracts textual descriptions of the coordinates from OpenStreetMap, which provides a sufficient geographic context. LLMs process these descriptions, and the final hidden states of individual tokens are averaged to form the LLMGeovec embedding for each coordinate.
LLMGeovec is not only the first geolocation representation model to achieve global coverage using LLMs, it also presents a simple yet effective paradigm for enhancing spatio-temporal learning with LLMs. In GP, LLMGeovec can be used either as a stand-alone representation or concatenated with existing representations generated using street view and human mobility data. In LTSF, different timing patterns of various nodes need to be handled, and the challenge is how to not only identify the unique characteristics of nodes but also model the correlations between them  (Nie et al. 2024a, 2022; Zeng et al. 2022). LLMGeovec can be concatenated to the temporal features of individual nodes, allowing the model to distinguish between different nodes while modeling their geographical connections. In terms of GSTF, current approaches focus on capturing spatial dependencies by using GNNs to aggregate node features with the guidance of an adjacency matrix (Shao et al. 2022a, d, c; Wu et al. 2019a; Chen et al. 2020; Geng et al. 2019). Adding LLMGeovec as new node features before performing GNNs can provide more a priori knowledge about the spatial semantics of nodes.
@@
  • Through extensive experimental analysis, we demonstrate that LLMGeovec achieves global coverage and significantly boosts the performance of leading GD, LTSF, and GSTF models.
@@
Currently, there are three primary types of geolocation representation models: GNN-based models, image-based models, and natural language processing (NLP)-based models. GNN-based models construct graphs from correlations between locations, such as geographic distance, points of interest (POI), and human mobility patterns. These models generate node representations through message passing on the constructed graphs (Zhang et al. 2021; Kim and Yoon 2022; Zhang et al. 2023a; Zhou et al. 2023b). In contrast, Image-based models utilize street view or satellite imagery and employ contrastive learning to generate representations tied to specific coordinates (Jean et al. 2019; Wang, Li, and Rajagopal 2020; Liu et al. 2023c; Li et al. 2022a; Xi et al. 2022). Meanwhile, NLP-based models leverage textual representations to represent the textual descriptions associated with the corresponding locations.
@@
Geolocation Representation Learning. Given a set of nodes $P\in\mathbb{R}^{2\times N}$, where $N$ represents the number of nodes, and $P_{i}=(P_{i}^{\text{lon}}$, $P_{i}^{\text{lat}})$ denotes the longitude and latitude of the $i$-th node. The goal of geolocation representation learning is to construct an effective encoder $f$ that transforms $P$ into geographically informative representations $Z=f(P)$. Here, $Z\in\mathbb{R}^{M\times N}$ with $M$ denoting the dimensionality of the representation.
GP. Given a set of nodes $P\in\mathbb{R}^{2\times N}$, each node is associated with geographic attributes such as climatic indicators (e.g., average annual temperature, humidity) and social indicators (e.g., regional average educational attainment, average annual income, poverty rate, crime rate). For a given set of geographic attributes $A\in\mathbb{R}^{1\times N}$, GP in the context of geolocation representation learning involves training a linear regressor with $Z_{[:K]}=f(P_{[:K]})$ to fit $A_{[:K]}$ using $K$ training samples. The performance of the regressor on the test sets $Z_{[K:]}=f(P_{[K:]})$ and $A_{[K:]}$ is used to measure the prediction task performance and the quality of the encoder $f$ and representations $Z$.
LTSF. We consider a historical multivariate time series (MTS) $X\in\mathbb{R}^{H\times N}$, where $N$ represents the number of nodes (variates) and $H$ is the number of historical time slots. The objective is to predict future values $Y\in\mathbb{R}^{F\times N}$, with $F$ as the number of future time slots. Each value of node $i$ at time slot $t$ is denoted by $X_{t}^{i}$, and their coordinates by $P\in\mathbb{R}^{2\times N}$.
@@
As depicted in Fig.  1, the proposed LLMGeovec method encapsulates two primary phases: prompt generation and text embedding via LLMs. Initially, we generate geographic descriptions based on specified coordinates leveraging map data. These descriptions are then transformed into embeddings by LLMs. The obtained embeddings can be used for GP, LTSF, and GSTF.
@@
  • Address: Utilizes reverse-geocoding to detail the hierarchy of location, from local neighborhoods to national identifiers.
@@
Consistent with many previous studies, high-quality geolocation representations can be used for GP with the help of partial region labeling (Wang et al. 2022; Jean et al. 2019; Lee et al. 2021; Wang, Li, and Rajagopal 2020; Zhang et al. 2021, 2023a; Zhou et al. 2023b; Kim and Yoon 2022). This is a direct application of LLMGeovec. To be specific, we will divide the locations into a training set and a test set, and use linear regression in the training set to map geolocation representation to location attributes. This is followed by testing in the test set. It is worth noting that since LLMGeovec achieves global coverage, it can be used in GPs of various scales (global, country, city) and can also be combined with other geolocation representations through feature concatenation.
@@
In this section, we describe the integration of LLMGeovec with LTSF models. We start by outlining a general LTSF model, which typically consists of a token embedding layer $E$, an encoder $C$, and a predictor $D$ (Chen et al. 2023; Li et al. 2023; Liu et al. 2023b; Yi et al. 2024). The embedding layer $E$ projects the the $t$-th historical record $X_{t}\in\mathbb{R}^{1\times N}$ into hidden temporal embeddings $S_{t}=E(X_{t})\in\mathbb{R}^{d_{t}\times N}$, where $d_{t}$ is the embedding dimension. Note that there can be a normalization operation in this embedder such as RevIN (Kim et al. 2021) to address the nonstationarity of time series. The encoder $C$ then models the node-to-node and slot-to-slot relationships across $H$ historical time slots, and the predictor $D$ generates predictions $\hat{Y}\in\mathbb{R}^{F\times N}$ for the future $F$ time slots. This process is formulated as follows:
@@
where the model parameters are updated automatically through gradient descent. In practice, the encoder $C$ can be instantiated by Transformer blocks, convolution, and MLPs, to model either channel dependencies or token correlations. Many state-of-the-art LTSF models follow this architectural template, such as TSMixer (Chen et al. 2023), RMLP (Li et al. 2023), iTransformer (Liu et al. 2023b), and FreTS (Yi et al. 2024). For other Transformer-based architectures that employ token-wise embedding, such as Autoformer (Wu et al. 2021) and Informer (Zhou et al. 2021), we can adapt them with a simple inverting strategy (Liu et al. 2023b).
@@
Previous spatio-temporal prediction models often employ GNNs to capture spatial relationships between nodes, aggregating them into node features, which are then input into the temporal modeling component sequentially or alternatively (Shao et al. 2022a, d, c; Wu et al. 2019a; Tang, He, and Zhao 2022; Tang et al. 2022). We refer to the spatio-temporal model as STGNN. Given the graph $G$, the structural template is framed as follows:
@@
Note that this scheme is applicable for STGNNs with different types of processing methods. Specifically, different models have different instantiations of Eq. 9, e.g., spatio-temporal message passing mechanisms. Both the mainstream time-then-space and time-and-space STGNN family discussed in Cini et al. (2023) can be seamlessly adopted simply by concatenating LLMGeovec into the input of each node.
@@
We study the effectiveness of LLMGeovec through extensive experiments. We first demonstrate that in geolocation representation models, LLMGeovec performs SoTA in GP tasks at all three scales: city, national, and global, and even outperforms end-to-end supervised training models. We examine two LLMs, LLaMa3 8B and Mistral 8x7B, both of which are able to produce high-quality geolocation representations with LLMGeovec. Moreover, LLMGeovec is seamlessly embedded into various LTSF and GSTF models and directly improves the model performance under various tasks. Notably, in the GSTF tasks, the use of a simple MLP and LLMGeovec outperforms many GNN-based approaches, and LLMGeovec shows great potential as an alternative to time-consuming GNNs. For a detailed description of the models and the datasets (GP, LTSF, GTSF), please refer to Appendix.
@@
A Multi-scale and Multi-topic GP Benchmark. As illustrated in Table 1, at the global scale, we collect 14 GP tasks. These include three climate indicators such as Annual Air Temperature and 11 social indicators like Population Density and Human Modification (detailed descriptions please refer to Appendix). We utilize 100,000 locations with global coverage, which are generated by  Manvi et al. (2024) (Africa: 19,855; Asia: 55,893; Europe: 6,825; North America: 8,440; South America: 5,189; Oceania: 2,049). In line with  Manvi et al. (2023, 2024), each GP task is associated with a corresponding GeoTIFF file. For each coordinate, the average value of 12 pixels surrounding the coordinate is taken as the value of the coordinate. Following the protocol of  Kim and Yoon (2022), we perform five cross-validation using ridge linear regression implemented in Sklearn  (Feurer et al. 2020; McDonald 2009), and the average of mean absolute error (MAE), root mean square error (RMSE) and $R^{2}$ are reported. On the country and city scales, we use existing benchmarks, including the social indicators in India  (Lee et al. 2021) and NYC  (Zhou et al. 2023b). We also employ ridge linear regression in Sklearn and report the MAE, RMSE, and $R^{2}$ on the test sets.
Baselines. We compare the LLMGeovec generated by LLaMa3 8B  (Touvron et al. 2023) and Mistral 8x7B  (Jiang et al. 2023). As shown in Tab.  2, we also compare the text embedding generated with Bert-whitening  (Su et al. 2021). In the city and country scales, we additionally compare Image-based and GNN-based geolocation representation models.
@@
Tasks LLMGeovec (LLaMa 3 8B) LLMGeovec (Mistral 8 x 7B) Bert-whitening (Bert base)
MAE RMSE R2 MAE RMSE R2 MAE RMSE R2
Annual Air Temperature 9.90 14.32 0.95 11.05 16.03 0.94 24.03 32.73 0.76
Annual Precipitation 2016.60 3021.76 0.86 2176.82 3245.12 0.83 3717.73 5293.82 0.56
Monthly Climate Moisture 1302.14 2021.21 0.55 1345.82 2097.32 0.52 1644.56 2715.44 0.19
Population Density 695.10 1020.11 0.85 759.81 1115.38 0.82 1342.98 2185.67 0.30
Nighttime Light Intensity 3.55 4.58 0.97 3.79 4.89 0.96 8.55 11.37 0.81
Human Modification Terrestrial 0.07 0.09 0.78 0.07 0.09 0.75 0.12 0.15 0.39
Global Gridded Relative Deprivation 6.56 8.98 0.85 6.70 9.15 0.84 10.43 13.60 0.65
Ratio of Built-up Area to Non-built Up Area 8.44 11.07 0.78 8.72 11.41 0.77 13.04 16.40 0.52
Child Dependency Ratio 5.84 8.29 0.86 5.88 8.34 0.88 9.50 13.12 0.64
Subnational Human Development 5.79 8.20 0.89 5.82 8.22 0.89 9.81 13.30 0.70
Infant Mortality Rates 3.98 6.06 0.93 4.02 6.14 0.93 7.42 10.76 0.77
Asset Index 0.02 0.03 0.93 0.02 0.03 0.92 0.06 0.08 0.53
Sanitation Index 0.09 0.12 0.95 0.10 0.13 0.93 0.23 0.30 0.67
Women BMI 0.76 1.01 0.95 0.83 1.12 0.94 1.82 2.38 0.77
@@
Datasets and models. Following the settings of  Zhang et al. (2024a); Wu et al. (2022); Shao et al. (2022a), we select five LTSF datasets from a wide range of domains, including Solar Energy, Global Wind, Global Temperature, Traffic flow, Delivery demand, and air quality. Several representative LTSF models are selected, including both Transformer-based and MLP-based methods. They are iTransformer (Liu et al. 2023b), TSMixer (Chen et al. 2023), RMLP (Li et al. 2023), and Informer (Zhou et al. 2021).
@@
Performances Comparisions using different geolocation embeddings. In this section, we compare the effects of two geolocation representations on LTSF models, and for reference, we also add learnable embeddings  (Shao et al. 2022b) (e.g., STID) with the same feature dimensions as LLMGeovec. For a fair comparison, all three methods use the same Adapter and model parameters. As shown in Tab.  7, LLMGeovec, which contains richer spatial semantics, achieves the greatest improvement.
@@
Hyperparameters Settings. We adapt the suggested hyperparameters in LargeST benchmark (Liu et al. 2023a) for all models.
@@
Performances in zero-shot scenarios. In addition to enhancing various models in full training scenarios for GSTF, LLMGeovec also has the potential for enhancing zero-shot transfer. We compare the performance of the learnable node embedding (STID) introduced by Shao et al. (2022a) and LLMGeovec in zero-shot scenarios. As a reference, we also test the transferability of the baseline GWNET and MLP. The LaDe data is adopted for this experiment. It is evident from Fig. 2 that when models are transferred to a new region in a zero-shot scenario, the learnable embedding harms the performance of MLP significantly because the embedding has adapted to the source data with specific patterns. In contrast, universal LLMGeovec can be generalized to other regions without any adjustment. This indicates that LLMGeovec features intrinsic geolocation knowledge that is generalizable for different regions.

### v2

@@ Abstract
In the geospatial domain, universal representation models are significantly less prevalent than their extensive use in natural language processing and computer vision. This discrepancy arises primarily from the high costs associated with the input of existing representation models, which often require street views and mobility data. To address this, we develop a novel, training-free method that leverages large language models (LLMs) and auxiliary map data from OpenStreetMap to derive geolocation representations (LLMGeovec). LLMGeovec can represent the geographic semantics of city, country, and global scales, which acts as a generic enhancer for spatio-temporal learning. Specifically, by direct feature concatenation, we introduce a simple yet effective paradigm for enhancing multiple spatio-temporal tasks including geographic prediction (GP), long-term time series forecasting (LTSF), and graph-based spatio-temporal forecasting (GSTF). LLMGeovec can seamlessly integrate into a wide spectrum of spatio-temporal learning models, providing immediate enhancements. Experimental results demonstrate that LLMGeovec achieves global coverage and significantly boosts the performance of leading GP, LTSF, and GSTF models. Our codes are available at https://github.com/Umaruchain/LLMGeovec.
@@ Introduction
The reason why previous approaches do not further extend the complex applications is that they do not achieve global coverage and they have a heavy reliance on expensive input data such as street views, travel patterns, and traffic trajectories  (Wang, Li, and Rajagopal 2020; Kim and Yoon 2022; Lee et al. 2021; Zhang et al. 2023a). Although studies have used free and globally available satellite imagery for geolocation representation, their effectiveness has been hampered by the low resolution and absence of important features such as activity patterns  (Manvi et al. 2023; Robinson, Hohman, and Dilkina 2017; Head et al. 2017; Jean et al. 2019; Elmustafa et al. 2022; Xi et al. 2022; Ma et al. 2023; Sun 2024).
@@
In this paper, we introduce a novel, training-free method that uses LLM and OpenStreetMap auxiliary map data to derive geolocation representations (LLMGeovec). As illustrated in Fig.  1, our approach first extracts textual descriptions of the coordinates from OpenStreetMap, which provides a sufficient geographic context. LLMs process these descriptions, and the final hidden states of individual tokens are averaged to form the LLMGeovec embedding for each coordinate.
LLMGeovec is not only the first geolocation representation model to achieve global coverage using LLMs, it also presents a simple yet effective paradigm for enhancing spatio-temporal learning with LLMs. In GP, LLMGeovec can be used either as a standalone representation or concatenated with existing representations generated using street view and human mobility data. In LTSF, different temporal patterns of various nodes need to be handled, and the challenge is how to not only identify the unique characteristics of nodes but also model the correlations between them  (Nie et al. 2024a, 2022; Zeng et al. 2022). LLMGeovec can be concatenated to the temporal features of individual nodes, allowing models to distinguish between different nodes while modeling their geographical connections. In terms of GSTF, current approaches focus on capturing spatial dependencies by using GNNs to aggregate node features with the guidance of an adjacency matrix (Shao et al. 2022a, d, c; Wu et al. 2019a; Chen et al. 2020; Geng et al. 2019; Nie et al. 2023). Adding LLMGeovec as new node features before performing GNNs can provide more a priori knowledge about the spatial semantics of nodes.
@@
  • Extensive experimental analysis demonstrates that LLMGeovec achieves global coverage and significantly boosts the performance of leading GD, LTSF, and GSTF models.
@@
Currently, there are three primary types of geolocation representation models: GNN-based models, image-based models, and natural language processing (NLP)-based models. GNN-based models construct graphs from correlations between locations, such as geographic distance, points of interest (POI), and human mobility patterns. These models generate node representations through message passing on the constructed graphs (Zhang et al. 2021; Kim and Yoon 2022; Zhang et al. 2023a; Zhou et al. 2023b). In contrast, Image-based models utilize street view or satellite imagery and employ contrastive learning to generate representations tied to specific coordinates (Jean et al. 2019; Wang, Li, and Rajagopal 2020; Liu et al. 2023c; Li et al. 2022a; Xi et al. 2022; Ma, Ni, and Chen 2024). Meanwhile, NLP-based models leverage textual representations to represent the textual descriptions associated with the corresponding locations.
@@
Geolocation Representation Learning. Given a set of nodes $P\in\mathbb{R}^{2\times N}$, where $N$ represents the number of nodes, and $P_{i}=(P_{i}^{\text{lon}}$, $P_{i}^{\text{lat}})$ denotes the longitude and latitude of the $i$-th node, the goal of geolocation representation learning is to construct an effective encoder $f$ that transforms $P$ into geographically informative representations $Z=f(P)\in\mathbb{R}^{M\times N}$ with $M$ denoting the dimension of the representation.
GP. Given a set of nodes $P\in\mathbb{R}^{2\times N}$, each node is associated with geographic attributes such as climatic indicators (e.g., average annual temperature, humidity) and social indicators (e.g., regional average educational attainment, average annual income, poverty rate, crime rate). For a given set of geographic attributes $A\in\mathbb{R}^{1\times N}$, GP in the context of geolocation representation learning involves training a linear regressor with $Z_{[:K]}=f(P_{[:K]})$ to fit $A_{[:K]}$ using $K$ training samples. The performance of the regressor on the test sets $Z_{[K:]}=f(P_{[K:]})$ and $A_{[K:]}$ is used to measure the quality of the encoder $f$ and the representations $Z$.
LTSF. We consider a multivariate time series (MTS) $X\in\mathbb{R}^{H\times N}$, where $N$ represents the number of nodes (variates) and $H$ is the number of historical time slots. The objective is to predict future values $Y\in\mathbb{R}^{F\times N}$, with $F$ as the number of future time slots. Each value of node $i$ in time slot $t$ is denoted by $X_{t}^{i}$, and their coordinates by $P\in\mathbb{R}^{2\times N}$.
@@
As depicted in Fig.  1, the proposed LLMGeovec encapsulates two primary phases: prompt generation and text embedding via LLMs. Initially, we generate geographic descriptions based on specified coordinates with map data. These descriptions are then transformed into embeddings by LLMs. The obtained embeddings can be used for GP, LTSF, and GSTF.
@@
  • Address: uses reverse-geocoding to detail the hierarchy of location, from local neighborhoods to national identifiers.
@@
Consistent with many previous studies, high-quality geolocation representations can be used for GP with the help of partial region labeling (Wang et al. 2022; Jean et al. 2019; Lee et al. 2021; Wang, Li, and Rajagopal 2020; Zhang et al. 2021, 2023a; Zhou et al. 2023b; Kim and Yoon 2022). This is a direct application of LLMGeovec. Specifically, we divide the locations into a training set and a test set, and use linear regression in the training set to map geolocation representation to location attributes. This is followed by testing in the test set. It is worth noting that since LLMGeovec achieves global coverage, it can be used in GPs of various scales (global, country, city) and can also be combined with other geolocation representations through feature concatenation.
@@
In this section, we describe the integration of LLMGeovec with LTSF models. We start by outlining a general LTSF model, which typically consists of a token embedding layer $E$, an encoder $C$, and a predictor $D$ (Chen et al. 2023; Li et al. 2023a; Liu et al. 2023b; Yi et al. 2024; Zhang, Guo, and Wang 2023). The embedding layer $E$ projects the the $t$-th historical record $X_{t}\in\mathbb{R}^{1\times N}$ into hidden temporal embeddings $S_{t}=E(X_{t})\in\mathbb{R}^{d_{t}\times N}$, where $d_{t}$ is the embedding dimension. Note that there can be a normalization operation in this embedder such as RevIN (Kim et al. 2021) to address the nonstationarity of time series. The encoder $C$ then models the node-to-node and slot-to-slot relationships across $H$ historical time slots, and the predictor $D$ generates predictions $\hat{Y}\in\mathbb{R}^{F\times N}$ for the future $F$ time slots. This process is formulated as follows:
@@
where the model parameters are updated automatically through gradient descent. In practice, the encoder $C$ can be instantiated by Transformer blocks, convolution, and MLPs, to model either channel dependencies or token correlations. Many state-of-the-art LTSF models follow this architectural template, such as TSMixer (Chen et al. 2023), RMLP (Li et al. 2023a), iTransformer (Liu et al. 2023b), and FreTS (Yi et al. 2024). For other Transformer-based models that employ token-wise embedding, such as models in (Wu et al. 2021; Zhou et al. 2021; Zhang, Wang, and Zhang 2024), we can adapt them with a simple inverting strategy (Liu et al. 2023b).
@@
Previous spatio-temporal prediction models often employ GNNs to capture spatial relationships between nodes, aggregating them into node features, which are then input into the temporal modeling component sequentially or alternatively (Shao et al. 2022a, d, c; Wu et al. 2019a; Tang, He, and Zhao 2022; Tang et al. 2022; Zhang et al. 2025). We refer to the spatio-temporal model as STGNN. Given the graph $G$, the structural template is framed as follows:
@@
Note that this scheme is applicable for STGNNs with different types of processing methods. Specifically, different models have different instantiations of Eq. 9, e.g., spatio-temporal message passing mechanisms. Both the mainstream time-then-space and time-and-space STGNN family discussed in (Cini et al. 2023) can be seamlessly adopted simply by concatenating LLMGeovec into the input of each node.
@@
We study the effectiveness of LLMGeovec through extensive experiments. We first demonstrate that in geolocation representation models, LLMGeovec performs SoTA in GP tasks at all three scales: city, national, and global, and even outperforms end-to-end supervised training models. We examine two LLMs, LLaMa3 8B and Mistral 8x7B, both of which are able to produce high-quality geolocation representations with LLMGeovec. Moreover, LLMGeovec is seamlessly embedded into various LTSF and GSTF models and directly improves the model performance under various tasks. Notably, in the GSTF tasks, the use of a simple MLP and LLMGeovec outperforms many GNN-based approaches, and LLMGeovec shows great potential as an alternative to time-consuming GNNs. For a detailed description of the models and the datasets (GP, LTSF, GSTF), please refer to Appendix.
@@
A Multi-scale and Multi-topic GP Benchmark. As illustrated in Table 1, at the global scale, we collect 14 GP tasks. These include three climate indicators such as Annual Air Temperature and 11 social indicators like Population Density and Human Modification (detailed descriptions please refer to Appendix). We use 100,000 locations with global coverage, generated by  Manvi et al. (2024) (Africa: 19,855; Asia: 55,893; Europe: 6,825; North America: 8,440; South America: 5,189; Oceania: 2,049). In line with  Manvi et al. (2023, 2024), each GP task is associated with a corresponding GeoTIFF file. For each coordinate, the average value of 12 pixels surrounding the coordinate is taken as the value of the coordinate. Following the protocol of  Kim and Yoon (2022), we perform five cross-validation using ridge linear regression implemented in Sklearn  (Feurer et al. 2020; McDonald 2009), and the average of mean absolute error (MAE), root mean square error (RMSE) and $R^{2}$ are reported. On the country and city scales, we use existing benchmarks, including social indicators in India  (Lee et al. 2021) and NYC  (Zhou et al. 2023b). We also employ ridge linear regression in Sklearn and report MAE, RMSE, and $R^{2}$ on test sets.
Baselines. We compare the LLMGeovec generated by LLaMa3 8B  (Touvron et al. 2023) and Mistral 8x7B  (Jiang et al. 2023). As shown in Tab.  2, we also compare the text embedding generated by Bert-whitening, GTE-large, and GTE-qwen2 7B  (Su et al. 2021; Li et al. 2023b). In the city and country scales, we additionally compare Image-based and GNN-based geolocation representation models.
@@
Tasks LLMGeovec (LLaMa 3 8B) LLMGeovec (Mistral 8 x 7B) Bert-whitening (Bert base) GTE-large GTE-qwen2 7B
MAE RMSE R2 MAE RMSE R2 MAE RMSE R2 MAE RMSE R2 MAE RMSE R2
Annual Air Temperature 9.90 14.32 0.95 11.05 16.03 0.94 24.03 32.73 0.76 22.23 30.15 0.80 14.05 19.99 0.91
Annual Precipitation 2016.60 3021.76 0.86 2176.82 3245.12 0.83 3717.73 5293.82 0.56 3519.57 5012.55 0.61 2604.86 3931.23 0.76
Monthly Climate Moisture 1302.14 2021.21 0.55 1345.82 2097.32 0.52 1644.56 2715.44 0.19 1619.65 2610.16 0.25 1394.10 2288.89 0.43
Population Density 695.10 1020.11 0.85 759.81 1115.38 0.82 1342.98 2185.67 0.30 1266.82 1986.15 0.42 896.52 1417.02 0.70
Nighttime Light Intensity 3.55 4.58 0.97 3.79 4.89 0.96 8.55 11.37 0.81 7.63 9.99 0.85 4.77 6.15 0.94
Human Modification Terrestrial 0.07 0.09 0.78 0.07 0.09 0.75 0.12 0.15 0.39 0.11 0.14 0.47 0.08 0.11 0.68
Global Gridded Relative Deprivation 6.56 8.98 0.85 6.70 9.15 0.84 10.43 13.60 0.65 9.83 12.95 0.68 8.13 10.86 0.78
Ratio of Built-up Area to Non-built Up Area 8.44 11.07 0.78 8.72 11.41 0.77 13.04 16.40 0.52 12.51 15.81 0.56 10.48 13.41 0.68
Child Dependency Ratio 5.84 8.29 0.86 5.88 8.34 0.88 9.50 13.12 0.64 9.11 12.47 0.68 7.17 10.09 0.79
Subnational Human Development 5.79 8.20 0.89 5.82 8.22 0.89 9.81 13.30 0.70 9.15 12.36 0.75 7.10 9.95 0.83
Infant Mortality Rates 3.98 6.06 0.93 4.02 6.14 0.93 7.42 10.76 0.77 7.24 10.20 0.80 4.97 7.50 0.89
Asset Index 0.02 0.03 0.93 0.02 0.03 0.92 0.06 0.08 0.53 0.05 0.07 0.62 0.04 0.06 0.78
Sanitation Index 0.09 0.12 0.95 0.10 0.13 0.93 0.23 0.30 0.67 0.20 0.26 0.75 0.15 0.20 0.85
Women BMI 0.76 1.01 0.95 0.83 1.12 0.94 1.82 2.38 0.77 1.55 2.04 0.83 1.16 1.57 0.90
@@
Performance of Prompt Variants. To examine the effect of the various parts of the prompts on the quality of LLMGeovec, we try several variants of the prompts and test them on GP tasks in NYC. As shown in Tab.  7, Address, which is the latitude and longitude counterpart of the location hierarchy, from local neighborhoods to national identifiers, is the most important part of motivating geographic knowledge in LLM. As for the K nearest places, we discover that when K is too small, it can not provide effective geographic information; when K is too large, the neighboring nodes will be too convergent, which leads to poor results in downstream tasks.
Table 7: Performance of Prompt Variants
Prompt Poverty Rate Education Level Income Level Crime Rate
MAE RMSE R2 MAE RMSE R2 MAE RMSE R2 MAE RMSE R2
Instruction + Address + Top 10 NearbyPlaces 0.13 0.74 0.62 0.39 0.17 0.75 0.61 0.36 0.12 0.65 0.53 0.34
Instruction + Address + Top 5 NearbyPlaces 0.17 0.75 0.61 0.36 0.12 0.65 0.53 0.34 0.06 0.61 0.48 0.32
Instruction + Address + Top 1 NearbyPlaces 0.12 0.65 0.53 0.34 0.06 0.61 0.48 0.32 -0.01 0.42 0.29 0.09
Instruction + Address 0.06 0.61 0.48 0.32 -0.01 0.42 0.29 0.09 0.13 0.74 0.62 0.39
Instruction -0.01 0.42 0.29 0.09 0.13 0.74 0.62 0.39 0.17 0.75 0.61 0.36
Performance of Regions with Sparse POIs. For the NearbyPlaces in the prompt, we use the node field in the OSM data, and when there are POIs (e.g., bar, restaurant) near the coordinates, NearbyPlaces will be a variety of such buildings or stores. If there are no POIs near the coordinates, NearbyPlaces will be the names of nearby streets. We show that LLMGeovec performs consistently in regions with extensive POI coverage (e.g., North America) and those with little coverage (e.g., Africa). Specifically, we utilize the continent boundaries to extract nodes belonging to North America and Africa separately and perform the five-fold cross-validation within the continents, the results of which are shown in Tab  8. where LLMGeovec’s results are similar in both regions.
Table 8: Performance of Regions with Sparse POIs.
Region Population Density Night Light Density Annual Air Temperature
Africa 0.90 0.94 0.88
North America 0.82 0.97 0.92
@@
Datasets and models. Following the settings of  Zhang et al. (2024a); Wu et al. (2022); Shao et al. (2022a), we select five LTSF datasets from a wide range of domains, including Solar Energy, Global Wind, Global Temperature, Traffic flow, Delivery demand, and air quality. Several representative LTSF models are selected, including both Transformer-based and MLP-based methods. They are iTransformer (Liu et al. 2023b), TSMixer (Chen et al. 2023), RMLP (Li et al. 2023a), and Informer (Zhou et al. 2021).
@@
Performances Comparisions using different geolocation embeddings. In this section, we compare the effects of two geolocation representations on LTSF models, and for reference, we also add learnable embeddings  (Shao et al. 2022b) (e.g., STID) with the same feature dimensions as LLMGeovec. For a fair comparison, all three methods use the same Adapter and model parameters. As shown in Tab.  9, LLMGeovec, which contains richer spatial semantics, achieves the greatest improvement.
@@
Hyperparameters Settings. We adopt the suggested hyperparameters in LargeST (Liu et al. 2023a) for all models.
@@
Overhead of LLMGeovec. For the GTSF and LTSF tasks, we concatenate the d-dimensional LLMGeovec to the original feature vector while reducing the dimensionality of the original feature vector by d. This way, the overall dimensionality remains unchanged. The only additional parameters introduced are the LLM Adapter (A two-layer MLP that uses Leaky_Relu in the middle), which is used to reduce the input dimensionality of the LLMGeovec. As shown in  11, our model introduces only an acceptable number of additional parameters and maintains computational efficiency comparable to the original model.
Table 11: Ovearhead of LLMGeovec.
Model Parameters Running Speed GPU Memory
STGCN w/o LLMGeovec 3482K 3.62 s/it 3.5 GB
STGCN w/ LLMGeovec 3351K 3.58 s/it 3.5 GB
Performances in zero-shot scenarios. In addition to enhancing various models in full training scenarios for GSTF, LLMGeovec also has the potential for enhancing zero-shot transfer. We compare the performance of the learnable node embedding (STID) introduced by Shao et al. (2022a) and LLMGeovec in zero-shot scenarios. As a reference, we also test the transferability of the baseline GWNET and MLP. The LaDe data is adopted for this experiment. It is evident from Fig. 2 that when models are transferred to a new region in a zero-shot scenario, the learnable embedding harms the performance of MLP significantly because the embedding has adapted to the source data with specific patterns. In contrast, universal LLMGeovec can be generalized to other regions without any adjustment. This indicates that LLMGeovec features intrinsic geolocation knowledge that is generalizable for different regions. In addition, more advanced techniques such as test-time adaptation in graphs can be applied to achieve better few-shot performance (Sun et al. 2024).